when the current age of the mother is added to the present age of the daughter is 40 years. What will be their total age after five years?
Answer by Alan3354(69443)

when the current age of the mother is added to the present age of the daughter is 40 years. What will be their total age after five years?
===============
M + D = 40
(M+5) + (D+5) = M+D + 10
= 50
-------------------------------
PS "present" and "current" are not needed.
----
The age of the mother added to the age of the daughter is 40 years.
What will be the sum of their ages after five years?
----
That's better.
I might be pedantic.